Β If π‘Ž = 4𝑑^(βˆ’ 3/2) , 𝑠 = 16 π‘€β„Žπ‘’π‘› 𝑑 = 4, π‘Žπ‘›π‘‘ 𝑠 = 25 π‘€β„Žπ‘’π‘› 𝑑 = 6, find the equation of motion 𝑠 = 𝑓(𝑑) and the velocity function 𝑣(𝑑).Β 


1
Expert's answer
2021-03-01T15:00:58-0500

"v(t)=\int a(t)dt= \int 4t^{-3/2}dt= \frac{-8}{\sqrt{t}} +C_1"

"s(t)= \int v(t)dt =\int( -8t^{-0.5}+C_1)dt =-16\sqrt{t} +C_1t + C_2"


16=-32+4C1+C2

C2=48+4C1


"25=-16 \sqrt{6}+ 6C1+48+4C1"

"10C_1=-23+16 \sqrt{6}"

"C_1=1.6 \sqrt{6}-2.3"

"C_2=38.8+6.4 \sqrt{6}"

Answer: "v(t)= \frac{-8}{\sqrt{t}}+1.6\sqrt{6}-2.3"

"s(t)=-16\sqrt{t}+t(1.6 \sqrt{6}-2.3) +38.8+6.4\sqrt{6}"